# CardioAI: A Multimodal AI-based System to Support Symptom Monitoring and Risk Detection of Cancer Treatment-Induced Cardiotoxicity

### v2

@@ 5.2.4. Concerns
While remote monitoring systems have the potential to improve patient care through real-time data collection, clinicians expressed ethical concerns around their practical implementation, especially around timely interventions, ambiguity in responsibility, and the lack of patient education.
@@
A further concern related to timely intervention is the ambiguity of responsibility when critical alerts are generated. Clinicians expressed that, even when alerts are issued promptly, the uncertainty around who is responsible for acting on them adds another layer of complexity. As P11 pointed out, “identifying who’s gonna take charge if there’s a red flag that comes out is a significant challenge.” P10 expanded on this concern by posing critical questions about how these alerts would be managed in practice, given the constraints of staffing and availability:
@@
P10 further stressed that patients need to be aware of the system’s limitations during off-hours, “Patients need to know that if they told their chatbot that they had chest pain at 9 pm on a Friday… no one’s gonna check the chatbot until Tuesday when they get back.
@@ 6. Discussion
In this discussion, we explore the broader implications of our findings. We begin by discussing the potential of our system in extending clinicians’ capabilities beyond clinical settings (Section  6.1). We then propose design considerations for AI-assisted proactive decision-making (Section  6.2). Finally, we highlighted the technical, ethical, and privacy challenges that arise from implementing these AI-based systems, with particular attention to infrastructure, patient reliance, and data security concerns (Section  6.3).
@@ 6.2. Design Considerations for AI-Assisted Proactive Decision-Making
From our findings, one of the most significant potentials of our multimodal AI-based system is facilitating the transition from reactive to proactive decision-making in cancer treatment-induced cardiotoxicity for clinicians. In current workflows, clinicians primarily respond to symptoms only after they have manifested or been reported by patients. This reactive approach can delay the detection of cardiotoxicity and increase the risk of more severe outcomes. Clinicians expressed that our system, which allows for continuous monitoring of physiological data and symptoms, paired with predictive risk scores, provides them with actionable insights in real-time. They highlighted the potential of our system to enhance their ability to intervene earlier, enabling them to make informed and proactive decisions before critical conditions arise. This shift aligns with broader trends in healthcare towards preventive care in cardiotoxicity management, as previously discussed by  (Wagner, 1998; Darkins et al., 2008), and our system effectively translates it into actionable clinical practice.
We also identify crucial design considerations that come with the integration of AI-assisted proactive decision-making through continuous monitoring. A central design consideration is its seamless integration into existing clinical workflows. Given that clinicians are already working in high-pressure environments with well-established routines, the introduction of these systems needs to be carefully aligned with current practices. Our system minimizes disruption by automating routine tasks, streamlining data collection, and delivering real-time insights directly within the clinical context. Another critical factor is the risk of information overload, a common challenge with continuous data streams. Our system mitigates this by filtering and prioritizing the most relevant data, ensuring that clinicians only receive actionable information.
This shift toward continuous monitoring and real-time decision support also necessitates rethinking traditional clinical workflows and responsibilities. Clinicians raised concerns about accountability in a continuously monitored environment, particularly regarding who would monitor the data and intervene when abnormalities are detected, particularly during non-working hours. These concerns introduce questions about the liability and workload of 24/7 monitoring, which requires a fundamental shift from the traditional episodic care model. The need for constant vigilance and potential interventions may challenge current healthcare delivery structures, and addressing these issues will be crucial to the successful implementation of AI-based systems in clinical practice.
@@ 6.3. Technical, Ethical and Privacy Concerns
Our study reveals several ethical and practical concerns that arise when integrating AI-based continuous monitoring systems into clinical practice. A critical concern is the need for robust infrastructure support, including reliable internet connections, functioning devices, and interoperability between systems. Clinicians noted that technical failures, such as malfunctioning devices or poor Wi-Fi connectivity, could disrupt real-time symptom monitoring, delay interventions and undermine the effectiveness of the system. Ensuring reliable IT infrastructure and technical support will be critical for the success of these systems. In addition to infrastructure, human resources are essential for ongoing monitoring and maintenance. While AI can automate data collection and analysis, human oversight remains crucial for interpreting alerts and responding to abnormal data. Clinicians expressed concerns about the increased workload that this can place on healthcare teams, particularly in 24/7 monitoring environments. Managing these demands will require careful planning to avoid overburdening staff and ensure continuous, high-quality care.
Beyond technical challenges, there are ethical concerns. One key issue is the overreliance of patients on AI-based systems that can delay critical medical attention. Clinicians expressed concern that patients might assume logging symptoms into a system guarantees immediate intervention, even when healthcare teams are unavailable, without fully understanding the limitations of technology. This risk is further amplified in underserved communities where access to clinicians is already limited, potentially leading to a false sense of security.
@@
Privacy and security concerns are also central to the implementation of AI and LLM-based systems. As these technologies handle sensitive patient data, ensuring the confidentiality and integrity of this information becomes crucial. Our study revealed that clinicians are particularly concerned about how patient data is stored, transmitted, and accessed within AI-based systems. The increasing reliance on AI raises questions about data breaches and unauthorized access, which could undermine patient trust and pose significant legal risks. Addressing these concerns will require robust data governance frameworks that prioritize patient privacy while allowing for the effective use of AI in healthcare.
@@ 6.4. Future Directions & Limitations
Our work has several limitations. First, while we gathered valuable insights during the prototype design phase with input from eleven experts, our evaluation study included only four participants. Future studies should involve a more diverse group of participants and conduct broader evaluations. Second, our system was developed as a prototype and used as a design probe to collect feedback from clinicians during the evaluation. This approach may affect the validity and generalizability of our findings, as the results might differ if the system were deployed in real-world clinical settings.
@@
Our findings suggest several avenues for future research and system refinement. First, further research should focus on enhancing the usability and integration of AI-based systems within clinical workflows, ensuring that the interface, data presentation, and interaction design align with clinicians’ needs. This includes developing seamless integration with EHR systems to reduce friction in day-to-day clinical use. Second, fostering trust in AI tools is crucial, and future research should explore improving the transparency and interpretability of AI-generated outputs. This can be achieved through the implementation of explainable AI (XAI) technologies and by providing clinicians with clear, understandable explanations of risk predictions and decisions made by the system. Third, while this study focused on cardiotoxicity, there is a broader need to investigate how similar AI-based systems can be applied to other clinical areas where early detection and intervention are critical. By exploring the deployment of proactive AI tools in various healthcare contexts, future work can help improve patient outcomes and optimize decision-making processes across different conditions.
@@ 7. Conclusion
In this study, we explore the challenges that clinicians face in managing cancer treatment-induced cardiotoxicity and developed a multimodal AI-based system, CardioAI, to support symptom monitoring and risk prediction. Through participatory design sessions with 11 clinicians, we uncovered the complexity of managing cardiotoxicity and identified gaps in existing monitoring approaches. Our system provides continuous monitoring of symptoms via wearable devices and LLM-based VA, and explainable AI-based predictive risk scores, which offer actionable insights to support clinical decision-making. Evaluation by clinical experts highlighted the system’s ability to reduce information overload, streamline workflows, and support proactive decision-making. These findings contribute to the growing field of HCI in healthcare, providing a foundation for future development of technologies to address similar challenges in clinical decision-making tasks. We envision that our work can inspire future designs of multimodal AI-based systems.

### v3

@@ 5.2.4. Concerns
While remote monitoring systems have the potential to improve patient care through real-time data collection, clinicians expressed ethical concerns around their practical implementation, especially regarding timely interventions, ambiguity in responsibility, and the lack of patient education.
@@
A further concern related to timely intervention is the ambiguity of responsibility when critical alerts are generated. Clinicians expressed that, even when alerts are issued promptly, the uncertainty around who is responsible for acting on them adds another layer of complexity. As P11 pointed out, “identifying who’s gonna take charge if there’s a red flag that comes out is a significant challenge.” P7 expanded on this concern by posing critical questions about how these alerts would be managed in practice, given the constraints of staffing and availability:
@@
P7 further stressed that patients need to be aware of the system’s limitations during off-hours, “Patients need to know that if they told their chatbot that they had chest pain at 9 pm on a Friday… no one’s gonna check the chatbot until Tuesday when they get back.
@@ 6. Discussion
In this discussion, we explore the broader implications of our findings. We begin by discussing the potential of our system in extending clinicians’ capabilities beyond clinical settings (Section  6.1). We then propose design considerations for AI-assisted proactive decision-making (Section  6.2). We highlighted the technical, ethical, and privacy challenges that arise from implementing these AI-based systems, with particular attention to infrastructure, patient reliance, and data security concerns (Section  6.3). Finally, we presented the limitations and future directions of our work (Section 6.4).
@@ 6.2. Design Considerations for AI-Assisted Proactive Decision-Making
Our studies highlight the potential of CardioAI in supporting clinical decision-making for cancer treatment-induced cardiotoxicity. Clinicians emphasized that integrating continuous monitoring with predictive risk scores facilitates a proactive approach to managing cardiotoxicity, enabling earlier interventions and more informed decisions before critical conditions arise. This represents a significant shift from traditional reactive workflows, aligning with broader trends in healthcare toward preventive care, as discussed in prior work (Wagner, 1998; Darkins et al., 2008), and demonstrates how our system translates this approach into actionable clinical practice.
Additionally, we propose several design considerations critical for the successful deployment of our system in real-world settings. These insights also serve as valuable guidance for future designers developing multimodal AI-based systems for clinical applications.
6.2.1. Fostering Accurate User Expectations through UI and XAI Design.
A foundational design consideration for multimodal AI-based systems in clinical settings is fostering accurate user expectations by leveraging UI and explainable AI (XAI) design. Given the complexity of these systems and their deployment in high-stakes clinical environments, it is essential to communicate their roles, capabilities, and limitations to users. We suggest that such communication be seamlessly embedded within the system’s design to ensure accessibility and usability while mitigating risks of over-reliance or inappropriate use.
A key aspect of fostering accurate user expectations is to provide explicit delineation of the system’s roles, capabilities, and limitations. Misunderstandings about what the system can and cannot do can lead to misuse with potentially severe consequences. For example, clinicians in our study emphasized the importance of ensuring that patients understand the LLM-VA’s intended role as a tool for symptom tracking and not as a substitute for emergency care. Misuse in such situations could lead to delays in critical interventions, which may pose significant risks to patient safety. We suggest that designers use UI and XAI design strategies to integrate this information directly into the system, instead of relying on methods like training sessions, manuals, and tutorials, which may not always be practical or accessible. For instance, designers can embed tooltips within the interface to provide contextualized definitions and explanations of the system functionality and limitations and use visual cues, such as color-coded indicators, to highlight critical outputs. For example, in our system, AI-based predictive risk scores were accompanied by detailed definitions, and red color was used to signify the detection of red-flag symptoms to alert clinicians about high-risk scenarios.
In addition to defining system roles and limitations, we recommend that designers provide interpretable and actionable explanations of the system’s outputs. Clinicians in our study highlighted the need to understand how AI-based risk predictions are generated, what they signify, and how they can inform actionable insights. In response to this need, our system incorporates explainable risk scores and feature importance indicators, which outline the key factors influencing each prediction. Additionally, these risk scores are accompanied by actionable recommendations tailored to different risk levels, providing clinicians with clear guidance on potential next steps. We suggest that designers consider how to provide explanations tailored to the nature of the technologies used in the system, such as detailing the sources of outputs, clarifying the reasoning behind predictions, and presenting insights in a concise and actionable format.
Explainability and expectation management should also remain dynamic, evolving alongside technological advancements and clinical workflow changes. For example, when new features are added or algorithms are improved, interfaces should be updated to reflect these changes accordingly. This could better support ongoing dialogue between stakeholders and designers, fostering a collaborative relationship that builds trust and accountability over time.
6.2.2. Balancing Clinical Benefits and Added Workload.
Another design consideration for multimodal AI-based systems in clinical decision-making is ensuring that the systems deliver clinical benefits without imposing additional workload and cognitive burden on clinicians. While the potential of such systems lies in their ability to provide comprehensive information, their effectiveness hinges on how this information is curated, prioritized, and presented. Poorly designed systems could overwhelm users with excessive information, reduce efficiency, and hinder clinicians’ ability to make timely and informed decisions.
To achieve additional clinical benefits, such systems often require the collection of extra data or the introduction of new information into clinicians’ workflows. For example, in our study, clinicians highlighted the tension between the value of having more data and the potential for information overload. Integrating AI-generated risk scores into the system provided valuable insights but also required clinicians to spend additional time interpreting these scores, assessing their accuracy, and evaluating their relevance. This inherent trade-off between added benefits and increased cognitive effort presents a significant design challenge. To address this, we recommend that designers leverage implicit user interactions with newly introduced information or functionalities as feedback to improve the system, taking this as an opportunity to minimize the additional workload introduced. This approach aligns with AI-in-the-loop principles, enabling systems to adapt dynamically to user behaviors. For instance, clinicians’ frequent engagement with specific flagged metrics could guide the system to refine its prioritization algorithms, thereby reducing cognitive effort and enhancing efficiency over time.
Another key aspect is designing systems that align with the diverse workflows and priorities of different clinical specialties. For instance, cardiologists in our study prioritized EKG data as a critical diagnostic tool, while oncologists placed greater emphasis on treatment plans and patient histories. This divergence underscores the importance of customizable system features, such as tailored notifications, adaptable dashboards, and role-specific thresholds. We recommend that designers actively collaborate with clinicians during the design process to identify these specialty-specific needs and ensure that systems accommodate them effectively. However, customization must be balanced with consistency: while personalized features can enhance relevance, preserving core functionalities across all user groups is essential to maintaining system usability, reliability, and safety.
@@ 6.3. Technical, Ethical and Privacy Concerns
While the scope of our study focuses on designing an interactive UI for the clinician-facing information dashboard to support clinicians’ decision-making, which does not extend to patient-facing interfaces or technologies, clinicians highlighted several concerns regarding the multimodal AI-based system for clinical decision-making. These concerns encompass infrastructure challenges, ethical considerations, and privacy risks which need to be addressed before real-world deployment.
6.3.1. Infrastructure Challenges
A critical concern is the need for robust infrastructure support, including reliable internet connections, functioning devices, and interoperability between systems. Clinicians noted that technical failures, such as malfunctioning devices or poor Wi-Fi connectivity, could disrupt real-time symptom monitoring, delay interventions, and undermine the effectiveness of the system. Ensuring reliable IT infrastructure and technical support will be critical for the success of these systems. In addition to infrastructure, human resources are essential for ongoing monitoring and maintenance. While AI can automate data collection and analysis, human oversight remains crucial for interpreting alerts and responding to abnormal data. Clinicians expressed concerns about the increased workload that this can place on healthcare teams, particularly in 24/7 monitoring environments. Managing these demands will require careful planning to avoid overburdening staff and ensure continuous, high-quality care.
6.3.2. Ethical Concerns
Beyond technical and resource challenges, clinicians raised ethical concerns, particularly around the potential for patients to over-rely on AI-based systems. Clinicians expressed concern that patients might assume logging symptoms into a system guarantees immediate intervention, even when healthcare teams are unavailable, without fully understanding the limitations of technology. This risk is further amplified in underserved communities where access to clinicians is already limited, potentially leading to a false sense of security.
@@ 6.3.2. Ethical Concerns
6.3.3. Privacy and Security Cocerns
Privacy and security concerns are also central to the implementation of AI and LLM-based systems. As these technologies handle sensitive patient data, ensuring the confidentiality and integrity of this information becomes crucial. Our study revealed that clinicians are particularly concerned about how patient data is stored, transmitted, and accessed within AI-based systems. The increasing reliance on AI raises questions about data breaches and unauthorized access, which could undermine patient trust and pose significant legal risks. Clinicians also highlighted the need for transparency in data handling practices to build trust among stakeholders. This includes clear communication about who has access to the data, how it is used, and the safeguards in place to protect it. Robust data governance frameworks are essential to address these concerns, ensuring compliance with legal and ethical standards while maintaining the confidentiality and integrity of patient information. While this study focuses on designing the clinician-facing dashboard for decision support, addressing privacy and security concerns across the entire multi-modal AI-based system – including patient-facing components such as wearables and voice assistants – will be critical. These aspects fall outside the scope of the current work but will be addressed in future efforts to design and integrate a complete remote monitoring and decision-making support system.
@@ 6.4. Future Directions & Limitations
Our work has several limitations. First, our sample sizes are limited. We involved 11 clinicians during the participatory design phase and four clinicians for the heuristic evaluation, similar to prior studies using small, expert-focused samples to gather domain-specific insights (Yang et al., 2024b; Zhang et al., 2024; Cai et al., 2019; Jacobs et al., 2021; Beede et al., 2020). Recruiting clinicians, especially in specialized domains like cardiotoxicity, presents unique challenges due to their demanding schedules, limited availability, and the niche nature of the expertise required. Despite these constraints, we considered the concept of information power, which posits that smaller, targeted samples can be sufficient when participants are highly specific to the study’s aims and the data collected is rich and relevant (Malterud et al., 2016). The recurring themes and consistent patterns observed across both phases further support the adequacy of our sample sizes in achieving the study’s objectives. However, all participants were recruited from a single hospital, which may introduce systematic biases. Also, our participant pool only included cardiologists and oncologists – the primary stakeholders in cardiotoxicity decision-making – other specialists, such as rheumatologists, radiologists, and nurses, could offer valuable perspectives depending on the specific cancer types or treatment pathways. We suggest future work could expand the participant pool to include a more diverse range of roles and institutions to enhance the generalizability of findings further and explore the system’s applicability in different clinical contexts.
Second, our system is a proof-of-concept prototype and has not been integrated into EHR systems or deployed in real-world clinical settings. The current prototype was used as a design probe to elicit clinician feedback on usability and decision-supporting functionalities, serving as a preliminary feasibility evaluation in a simulated environment. By conducting this initial evaluation, we aimed to avoid prematurely investing significant time and effort from domain experts in a large-scale controlled usability experiment before addressing key design and functional concerns. While this approach provides valuable early insights, it may influence the validity and generalizability of our findings. As discussed in Section 6.3, real-world deployment may reveal additional challenges, such as seamless EHR integration, operational scalability, and privacy and ethical concerns. Future work should consider these challenges to ensure the system’s robustness and practical utility.
Third, our study specifically focused on the interactive UI design of the clinician-facing dashboard and its core functionalities from the clinicians’ perspectives. We explored how the clinician-facing dashboard could support clinicians’ decision-making in managing cancer treatment-induced cardiotoxicity. Evaluations of patient-facing components, such as patients’ interactions with the LLM-VA or their experiences using wearables for symptom tracking, were beyond the scope of this study. Future work should include comprehensive and systematic evaluations of these patient-facing modules, such as assessing the LLM-VA for symptom reporting and wearables for continuous monitoring. These aspects are critical for scaling the system to real-world deployments and ensuring it effectively meets the needs of all stakeholders, including both patients and clinicians.
Also, fostering trust in AI-based tools is fundamental, particularly in high-stakes clinical settings. In this study, we explored the potential of explainable AI technologies to help domain experts better understand the outputs of AI and LLM-based systems. However, trust is a multifaceted concept, encompassing not only technical transparency but also reliability, openness, tangibility, etc. Future work could delve deeper into strategies for fostering trust in LLM-VA and AI models, such as incorporating uncertainty metrics that highlight the confidence level of predictions, and explore the interplay between these strategies and users’ perceptions of trust.
@@
Last but not least, while this study focused on cardiotoxicity, there is a broader need to investigate how similar AI-based systems can be applied to other clinical areas where early detection and intervention are critical. By exploring the deployment of proactive AI tools in various healthcare contexts, future work can help improve patient outcomes and optimize decision-making processes across different conditions.
@@ 7. Conclusion
In this study, we explore clinicians’ challenges in managing cancer treatment-induced cardiotoxicity and develop a multimodal AI-based system, CardioAI, to support symptom monitoring and risk prediction. Through participatory design sessions with 11 clinicians, we uncovered the complexity of managing cardiotoxicity and identified gaps in existing monitoring approaches. Our system provides continuous monitoring of symptoms via wearable devices and LLM-based VA, and explainable AI-based predictive risk scores, which offer actionable insights to support clinical decision-making. Evaluation by clinical experts highlighted the system’s ability to reduce information overload, streamline workflows, and support proactive decision-making. These findings contribute to the growing field of HCI in healthcare, providing a foundation for future development of technologies to address similar challenges in clinical decision-making tasks. We envision that our work can inspire future designs of multimodal AI-based systems.
Acknowledgements.
This work was funded in part by the National Science Foundation under award number IIS-2145625, and IIS-2302730, by the National Institutes of Health under award number R01AI188576 and R01MD018424, and by The Ohio State University President’s Research Excellence Accelerator Grant and by The Northeastern University Tier-1 Research Grant. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Science Foundation or the National Institutes of Health. The authors extend heartfelt thanks to the participants from The Ohio State University Wexner Medical Center.
Appendix A Participatory Design Exit Interview Script
  • Question 1 - Background: Could you please tell me a bit more about yourself and your practice, such as your years of study, residence, daily workload, etc.?
  • Question 2 - Experience of Cancer Treatment-Induced Cardiotoxicity: Can you recall a recent cardiotoxicity encounter for a cancer patient during treatment or near the completion stage? When did it happen? What types of cancer and cancer treatment were given?
  • Question 3 - Diagnosis and Monitoring of Cancer Treatment-Induced Cardiotoxicity: When do you suspect cardiotoxicity? What symptoms or signals lead to your suspicion? How do you solve the problem?
  • Question 4 - Technology Uses and Potentials in Current Cancer Treatment-Induced Cardiotoxicity Management: What kind of technology have you been using? Are you aware of any technologies or tools that others in your field or institutions have been exploring or adopting for cancer treatments and cardiotoxicity management?
  • Question 5 - Perspectives Towards Potential Technologies: Are you familiar with or have you explored the potential uses of emerging technologies, such as wearables, chatbots, or AI-based risk prediction models? Do you think any of these technologies could benefit risk management and early diagnosis? Do you have any concerns? [What information do you wish AI could have provided you?][What metrics or symptoms do you think might be important to collect when a patient is at home?]
  • Question 6 - Closing Questions: Is there anything else that you would like to share with us or any questions you have for us?
Appendix B Heuristic Evaluation Exit Interview Script
  • Question 1 - Overall Feedback: Can you share your overall feedback on the system?
  • Question 2 - Information Use: Do you find the information provided helpful for assisting in decision-making? Was the amount of information appropriate (e.g., too much, too little, or just right)?
  • Question 3 - Interaction Design: How would you describe your overall experience with the system’s user interface and interaction design? What about data visualizations? [further ask in terms of intuitive design, information display, customization, navigation, and alerts]
  • Question 4 - Concerns: Do you have any concerns about using this system in clinical workflows?
  • Question 5 - Future Design: What improvements or additional features would you suggest for future iterations of the system? [e.g., Is there any additional information you wish had been included?]
Table 2. Qualitative Codebook of Participatory Design Study Findings (Part 1)
Theme Sub-Theme Example
Challenges in Cardiotoxicity Decision-making Symptoms Not Just Subtle and Infrequent, But Sometimes Absent in the Early Stages ”I got this message he had a severe form of arrhythmia… He was completely asymptomatic. He was sitting at home relaxing, and the monitor picked up this thing.” (P4)
”So it all depends. If the patient, symptomatic or not, if they’re asymptomatic, then we’re the ones picking it up.” (P6)
”If a toxicity arises, especially that they’re asymptomatic from, they’re very hesitant to want to reach out..” (P1)
Logistical Barriers Compound Self-Reporting Limitations ”I think as much as we encourage them to reach out to us if they’re having symptoms, many of them still feel like they’re bothering us if they call.” (P2)
”In some areas, patients may not hear back right away from their physicians or nurses if they’re having symptoms… They may feel like they’re complaining too much or being a burden.” (P1)
Non-Specific Tools Compound Self-Reporting Limitations ”It was it kind of presented itself (cardiotoxicity-related symptoms), and it had been almost 4 or 5 days before when she had first described the symptom.” (P2)
”We use a survey tool to ask patients how they’re feeling, but it’s not specific to cardiotoxicity..” (P5)
Absence of Risk Patterns as an Additional Burden to Clinician Workload ”What type of features correlate with the likelihood somebody’s going to have cardiotoxicity is still, I think, very shaky. From what I understand, I think we don’t have a good grasp.” (P1)
”Because he had no cardiac risk factors. He had no heart risk factors that could predict this side effect.” (P4)
Opportunities Continuous Monitoring of Key Clinical Metrics Using Wearables ”I think we can pick up more patients with this. Or we can pick up patients earlier with this… We may be able to find more cardiac patients or cardiac adverse events related to our medications by doing this kind of monitoring.” (P4)
”I don’t know what they are like outside of the clinic … Those things can sometimes be helpful. It’s almost like people who have cardiac monitors.” (P1)
Remote Monitoring for Geographically Distant Patients ”A lot of our patients… travel very long distances to come here for treatment… they may call and be having an event where they feel really lightheaded or like they almost passed out… Having that data, especially if they can’t come in right away, is really helpful.” (P1)
”Doing more for patients in rural areas could be a place where this could work pretty well.” (P11)
LLM-Based Conversational Agent for Symptom Tracking ”I asked them to keep checking at home and write it down on paper, and they bring it to me… But sometimes they forget, or you know things happen… the concept is great, and having the patient have control over the symptoms and speaking and just translated that into something that we can see.” (P3)
”This is a really good idea. And that way they talk, and they don’t need to be typing or spending time.” (P2)
Table 3. Qualitative Codebook of Participatory Design Study Findings (Part 2)
Theme Sub-Theme Example
Opportunities AI-based Summarization ”There’s a lot of burnout in medicine, and the burnout is from all of this documentation that we do. I see a huge role of AI in that documentation… fulfill out the EHR data automatically.” (P1)
”If you could summarize the medical record, I think that’d be really helpful. Instead of going all over the place, looking around, it would be helpful.” (P2)
AI-based Risk Prediction ”If you are able to do it upfront, it would be very helpful… you will have to talk to some cardiologists to see how you define the score.” (P4)
”I think with AI there will be tools in place to even suggest what the next order should be… AI would help you understand what the next step could be even before you start talking to the patient.” (P7)
”I have not come across any risk scoring for cardiotoxicity. So I think that would be helpful.” (P5)
Design Suggestions Monitoring Key Clinical Metrics ”Oxygen levels would be nice. I know a lot of these wearable devices can do that.” (P2)
”It’s heart rate. It’s basically heart rate. Yeah, that can be monitored.” (P4)
Monitoring Key Clinical Symptoms ”Chest discomfort is good, shortness of breath, palpitations… These are all symptoms that we are looking out for… If they’re symptomatic, they should be checked quickly.” (P4)
”This is good; shortness of breath, palpitation, especially palpitations.” (P8)
Patient-Specific Baseline Alerts for Remote Monitoring ”But there are normal ranges for all of these things, but I think the change is based on what the patient’s baseline is. There’s something called clinically meaningful change. That really depends on where the patient baseline is.” (P6)
”I think it’s most helpful to have a baseline because people come in with all variety of kind of where they were at before.” (P1)
Enhanced Visualizations for Critical Data Interpretation ”The trend is extremely more important than just the one time.” (P5)
”Whenever the patient comes for their infusions, whenever they have the encounters with their physician. That could also be a good time to review these trends and alerts and see how they would change treatment decisions.” (P8)
”I think that would be helpful if patients are having an event where they felt like they were gonna pass out, or they were having palpitations to them, being able to correlate the time with whatever their biometrics were at the same time can be helpful.” (P1)
Clear Explanation of Technology to Prevent Misunderstanding ”So you have to see what you build it (AI risk score) on… How will you define the score? The problem with scores is that you have to sit down and manually do that. If something does it for us, we’ll be happy to use it. … AI is supposed to be my assistant, not my replacement.” (P4)
”Whenever we get this data and these scores, what do we do with it? ..especially patients are going to see cause. At the end of the day, it’s just having somewhat clear guidelines on what score means.” (P1)